At a distance of 5 m from a point sound source, the sound intensity level is 110 dB. At what distance is the intensity level 95dB?
1
Expert's answer
2020-12-28T08:42:02-0500

The intencity of the sound produced by a point source is inversely proportional to the squared distance from this source. Thus, obtain:


"I_1\propto\dfrac{1}{r_1^2}\\
I_2\propto\dfrac{1}{r_2^2}"

where "I_1" and "I_2" are intencities at the distances "r_1 = 5m" and "r_2" respectively. Thus, obtain:


"r_2 = r_1\sqrt{\dfrac{I_1}{I_2}}"

By definition, dB are:


"110dB = 10\lg I_1\\
90dB = 10\lg I_2"

From this obtain:


"I_1 = 10^{11}\space (relative\space units)\\
I_2 = 10^{9}\space (relativ\space units)"

Exact units don't matter since we are only interested in ratio "I_1\/I_2". Substituting into the expression for "r_2", obtain:


"r_2 = 5\sqrt{\dfrac{10^{11}}{10^9}} = 50m"

Answer. 50m.